Determine the number of atoms in 3.50 mol Fe.


1
Expert's answer
2022-02-11T10:45:03-0500


Avogadro's Constant =

6.02214086 × 10^23 mol-1= 1mole

Atoms= 3.5×6.02214086 × 10^23


Atoms= 2.1077× 10^24